# SAGI: Semantically Aligned and Uncertainty Guided AI Image Inpainting

### v2

@@ 8.4 Qualitative Analysis of Localization Performance
9 Qualitative Analysis
Figures 10 and 12 present a qualitative analysis of some cases from our dataset. In Figure 10, we show successful inpainting examples across different models and datasets (COCO, RAISE, and OpenImages), where the models correctly follow the prompts while producing realistic results.
Figure 12 illustrates different types of failure modes. The top two rows reveal problems with LLM-generated prompts, showing cases where prompts either fail to match the scene context or lead to technically sound but unrealistic results. Row 3 demonstrates technical limitations with visible artifacts and blurs. Row 4 presents cases where the inpainting appears realistic but deviates from the given prompt. Row 5 shows examples of poor inpainting quality where the models fail to generate coherent content. Finally, row 6 illustrates a subtle failure mode where the inpainting is technically well-executed but produces results that appear unnatural to human observers upon closer inspection. While these cases might be easily identifiable as manipulated by careful observers, they could potentially deceive viewers who are not actively looking for signs of manipulation, highlighting the importance of including such examples in inpainting datasets for developing robust detection methods.
(a) ”a juicy orange to add a vibrant pop of color to the composition”
(b) ”a majestic snow-capped mountain to create a scenic landscape”
(c) ”a vibrant blue poison dart frog”
(d) ”a cozy blanket and fluffy pillows to complete the bedroom scene”
(e) ”a grand marble fountain surrounded by lush greenery”
(f) ”a decorative ceramic vase”
(g) ”a playful otter swimming in the river stream”
(h) ”a cluster of small red berries growing in the grass”
(i) ”a rustic wooden barrel planter”
(j) ”a fresh, delicious sandwich to complete the meal”
(k) ”a clear blue sky to enhance the mountain landscape”
(l) ”a ripe golden delicious apple”
(m) ”a delicious cheeseburger to make the meal even more tempting”
(n) ”a lush green meadow, adding a touch of nature to the serene landscape”
(o) ”an intricately carved wooden eagle head”
@@ 9 Qualitative Analysis
(a) ”a person playing volleyball on the beach”
(b) ”a colorful hot air balloon floating in the sky”
(c) ”a red London phone booth”
(d) ”a blue bus traveling down the tracks”
(e) ”a playful panda bear imitating a martial arts move”
(f) ”a fluffy orange tabby cat with bright blue”
(g) ”vibrant red cherries to create a fruity collage effect”
(h) ”a majestic deer with large ant”
@@
(j) ”a group of young adults playing frisbee”
(k) ”a bright red mailbox to blend seamlessly into the park scene”
(l) ”a colorful bowl of fruit salad”
(m) ”a majestic lion standing proudly in the savanna”
(n) ”a large, majestic white husky standing”
(o) ”a majestic eagle perched on”
(p) ”a woman wearing a scarf and holding a bouquet of flowers”
(q) ”a bowl of fresh fruits”
(r) ”a friendly raccoon walking across a stone wall near trees”

### v3

@@ 8.4 Qualitative Analysis
9 Example Outputs
Figures 11 and 13 present a qualitative analysis of some cases from our dataset. In Figure 11, we show successful inpainting examples across different models and datasets (COCO, RAISE, and OpenImages), where the models correctly follow the prompts while producing realistic results.
Figure 13 presents different failure modes. The top two rows reveal problems with LLM-generated prompts, showing cases where prompts either fail to match the scene context or lead to technically sound but unrealistic results. Row 3 demonstrates technical limitations with visible artifacts and blurs. Row 4 presents cases where the inpainting appears realistic but deviates from the given prompt. Row 5 shows examples of poor inpainting quality where the models fail to generate coherent content. Finally, row 6 illustrates a subtle failure mode where the inpainting is technically well-executed but produces results that appear unnatural to human observers upon closer inspection. While these cases might be easily identifiable as manipulated by careful observers, they could potentially deceive viewers who are not actively looking for signs of manipulation, highlighting the importance of including such examples in inpainting datasets for developing robust detection methods.
(a) “a juicy orange to add a vibrant pop of color to the composition”
(b) “a majestic snow-capped mountain to create a scenic landscape”
(c) “a vibrant blue poison dart frog”
(d) “a cozy blanket and fluffy pillows to complete the bedroom scene”
(e) “a grand marble fountain surrounded by lush greenery”
(f) “a decorative ceramic vase”
(g) “a playful otter swimming in the river stream”
(h) “a cluster of small red berries growing in the grass”
(i) “a rustic wooden barrel planter”
(j) “a fresh, delicious sandwich to complete the meal”
(k) “a clear blue sky to enhance the mountain landscape”
(l) “a ripe golden delicious apple”
(m) “a delicious cheeseburger to make the meal even more tempting”
(n) “a lush green meadow, adding a touch of nature to the serene landscape”
(o) “an intricately carved wooden eagle head”
@@ 9 Example Outputs
(a) “a person playing volleyball on the beach”
(b) “a colorful hot air balloon floating in the sky”
(c) “a red London phone booth”
(d) “a blue bus traveling down the tracks”
(e) “a playful panda bear imitating a martial arts move”
(f) “a fluffy orange tabby cat with bright blue”
(g) “vibrant red cherries to create a fruity collage effect”
(h) “a majestic deer with large ant”
@@
(j) “a group of young adults playing frisbee”
(k) “a bright red mailbox to blend seamlessly into the park scene”
(l) “a colorful bowl of fruit salad”
(m) “a majestic lion standing proudly in the savanna”
(n) “a large, majestic white husky standing”
(o) “a majestic eagle perched on”
(p) “a woman wearing a scarf and holding a bouquet of flowers”
(q) “a bowl of fresh fruits”
(r) “a friendly raccoon walking across a stone wall near trees”